OpenStudy (anonymous):

divide and simplify: (y^3+4y/y^2-16) /(y^2+4y-12/y^2+2y-24)

OpenStudy (anonymous):

(y^3+4y/y^2-16) /(y^2+4y-12/y^2+2y-24)= (y^3+4y/y^2-16) * (y^2+2y-24/y^2+4y-12)= {y(y^2 +4) /(y+4)(y-4)} * {(y+6)(y-4)/(y+6)(y-2)}= y(y^2 +4)/(y+4)(y-4) * (y-4)/(y-2) = y(y^2 +4)/(y+4)(y-2)